OpenStudy (anonymous):

1/5 +1/5 1/8 +1/8+1/8 2/5+1/5

OpenStudy (anonymous):

\[\left\{\frac{2}{5},\frac{3}{8},\frac{3}{5}\right\} \]

OpenStudy (radar):

Addition of fractions. Note that each expression have common denominators,........so all we have to do is combine the numerators. 1/5 + 1/5 = 2/5 3/8 3/5 That is all there is to it.